OpenStudy (anonymous):

how do you simplify this:(x+1)(x+1)/2(x+1)

OpenStudy (anonymous):

Is the question \(\huge\frac{(x + 1)(x + 1)}{2(x + 1)}\)?

OpenStudy (anonymous):

1/2(x+1)^3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Do you see how there is an x + 1 in both the top and bottom?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Since there is one in both the top and bottom, they cancel out. \[\huge\frac{(x + 1)(x + 1)}{2(x + 1)} = \huge\frac{\cancel{(x + 1)}(x + 1)}{2\cancel{(x + 1)}} = ?\]

OpenStudy (anonymous):

okay so it will be x+1/2

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

okay thanks!:)

OpenStudy (anonymous):

np :)